In the laboratory you dissolve 22.1 g of chromium(III) acetate in a volumetric flask and add water to a total volume of 500 mL.

What is the molarity of the solution? _______ M.

What is the concentration of the chromium(III) cation? ______ M.

What is the concentration of the acetate anion?__________ M
1
Expert's answer
2019-03-14T04:00:55-0400

Dissociation of the salt:

Cr(CH3COO)3->Cr3++3CH3COO-

i) Find molarity of the solution:

n=m/M

n(Cr(CH3COO)3=22.1g/229.15g/mol=0.0964mol

c=0.0964mol/0.5L=0.193mol/L=0.193M

ii) According to equation 1mole of Cr(CH3COO)3 gives 1 mole of Cr3+, then n(Cr3+)=n(Cr(CH3COO)3)=0.0964mol

c(Cr3+)=0.0964mol/0.5L=0.193mol/L=0.193M

iii) According to equation 1mol of Cr(CH3COO)3 gives 3 moles of CH3COO-, then n(CH3COO-)=3*n(Cr(CH3COO)3)=3*0.0964=

0.289mol

c(CH3COO-)=0.289mol/0.5L=0.579mol/L=0.579M.